OpenStudy (anonymous):

check two of my 3 questions and help me solve for the third one pls

OpenStudy (anonymous):

What percent of the total responses is Italian? Round your answer to the nearest whole number percent. What percent difference is there between the Italian group and Ranch group? Round your answer to the nearest whole number percent. Which categories have a total of 30 responses?

OpenStudy (anonymous):

my answer for the first is 21 and two 23.8

OpenStudy (anonymous):

What percent of the total responses is Italian? Round your answer to the nearest whole number percent. Total Responses: 25 + 5 + 18 + 45 + 32 + 15 = 140 Italian: 15 15/140 = 0.107242857 11% italian.

OpenStudy (anonymous):

i divided 140/100 for some reason to solve thx

OpenStudy (anonymous):

ok for two i got 12 % difference number 3 dont get what its asking

OpenStudy (anonymous):

What percent difference is there between the Italian group and Ranch group? Round your answer to the nearest whole number percent. Italian = 11 Ranch = 23 23 = 11 = 12%

OpenStudy (anonymous):

i just wrte thatt

OpenStudy (anonymous):

Which categories have a total of 30 responses? vinaigrette = 5 french = 25 25 + 5 = 30